• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Steffen

    Log entry time 16:31:50 on October 14,1998

    Entry number 2886

    keyword=Night Update, Target FSD

    An update of everything that has gone on since 4pm:

    16:00 - 16:30 An access to check shower block #23 in the Hadron Arm. The signal
    looks good.
    16:30-18:00 Waiting for beam.
    18:00-19:35 Moller
    19:35-21:00 We wanted to turn the beam on to check the raster but had a problem.
    MCC reported a flag for the the target motion FSD. After fooling around
    with the motion, we called Ole' for his help. He said that the FSD was
    tied to the same circuit that the target position LED labeled "Target
    Switches" on the Motion Controller. So, one needs to be sure that green
    LED is on when the target is in place.
    The raster looked ok.
    21:00-21:40 Took polarized He3 Long. data at 0.666 GeV.
    21:40-22:00 MCC asked to have the beam for a few minutes.
    22:00-22:20 Took last He3 Long. data a 0.666 GeV Run#1789